# Fine-Grained Image Quality Assessment for Perceptual Image Restoration

### v3

@@ Abstract
Recent years have witnessed remarkable achievements in perceptual image restoration (IR), creating an urgent demand for accurate image quality assessment (IQA), which is essential for both performance comparison and algorithm optimization. Unfortunately, the existing IQA metrics exhibit inherent weakness for IR task, particularly when distinguishing fine-grained quality differences among restored images. To address this dilemma, we contribute the first-of-its-kind fine-grained image quality assessment dataset for image restoration, termed FGRestore, comprising 18,408 restored images across six common IR tasks. Beyond conventional scalar quality scores, FGRestore was also annotated with 30,886 fine-grained pairwise preferences. Based on FGRestore, a comprehensive benchmark was conducted on the existing IQA metrics, which reveal significant inconsistencies between score-based IQA evaluations and the fine-grained restoration quality. Motivated by these findings, we further propose FGResQ, a new IQA model specifically designed for image restoration, which features both coarse-grained score regression and fine-grained quality ranking. Extensive experiments and comparisons demonstrate that FGResQ significantly outperforms state-of-the-art IQA metrics. Codes and model weights have been released in https://sxfly99.github.io/FGResQ-Homepage.
@@ Introduction
Perceptual image restoration (IR) stands as a cornerstone in low-level computer vision, aiming to recover high-quality images from their degraded observations while maintaining perceptual fidelity (potlapalli2023promptir; luo2023controlling). Recent years have witnessed remarkable breakthroughs in this field, largely driven by the evolution of generative models. These advances have enabled IR algorithms to achieve unprecedented visual quality, creating an urgent demand for accurate image quality assessment (IQA) methods that can reliably evaluate and compare restored images. Such assessment capabilities are essential not only for performance benchmarking across different IR algorithms but also for guiding algorithmic optimization.
@@
Currently, IR algorithm evaluation still predominantly relies on traditional reference-based metrics such as PSNR and SSIM (SSIM), which assess restoration quality by measuring the similarity between the restored images and ground-truth images, even though their inconsistency with human perceptual judgment has been increasingly observed by the research community (PIPAL). Recent efforts have introduced no-reference IQA methods (xu2024towards; zhou2025unires), including CLIP-IQA (CLIP-IQA) and other learning-based approaches (MUSIQ; Q-align), which prove particularly valuable in real-world restoration scenarios where ground truth images are unavailable. However, a critical oversight persists: these methods have been primarily designed for image quality assessment, without specific consideration for the fine-grained evaluation requirements inherent in IR tasks.
@@
To investigate this question, we first conducted comprehensive computational analyses on established IQA datasets for image restoration (PIPAL; SISRSet). Our findings consistently reveal a significant performance gap: while existing IQA methods achieve reasonable results in coarse-grained quality assessment, they consistently underperform in fine-grained scenarios, with significant inconsistencies between predicted scores and human perceptual judgments. This observation highlights the inadequacy of current evaluation protocols for capturing the nuanced quality differences which are crucial in IR applications. Moreover, when quality differences between images become subtle, both human and IQA models encounter significant challenges in providing reliable absolute quality scores. In contrast, pairwise comparison(10168279; yang2024semantics) presents a more reliable alternative for fine-grained assessment, as humans demonstrate superior consistency in relative quality judgments compared to absolute scoring.
@@
To evaluate IR algorithms, numerous datasets have been proposed. These include datasets dedicated to single restoration types, such as the MDD13 (MDD13) dataset for deblurring, the IVC-Dehazing (IVC-Dehazing) and exBeDDE (exBeDDE) datasets for dehazing, the IVIPC-DQA (IVIPC_DQA) dataset for deraining, and the QADS (QADS), SISRSet (SISRSet), RealSRQ (RealSRQ) and SRIQA-Bench (chen2025toward) datasets for super-resolution. Additionally, there are datasets that encompass distortions from various algorithms, such as PIPAL (PIPAL), which includes distortion types like denoising, super-resolution, and super-resolution with post-denoising, KADID-10k (KADID-10k), which contains various image distortions including denoising degradations and DiffIQA(chen2025toward), which accounts for degradation artifacts induced by diffusion-based image enhancement. A brief comparison of FGRestore and previous IR evaluation datasets is illustrated in Table 1.
@@
Existing image restoration (IR) algorithms typically employ full-reference image quality assessment (FR-IQA) metrics for performance evaluation, such as the conventional PSNR and SSIM (SSIM), as well as deep learning-based metrics like LPIPS (LPIPS) and DISTS (DISTS). To overcome the flawed ”perfect reference” assumption in traditional FR-IQA, which is challenged by imaging system limitations and superior generative methods, A-FINE (chen2025toward) proposes a generalized model that adaptively assesses both the fidelity and naturalness of test images.
No-reference IQA (NR-IQA) methods have been proposed to assess image quality without reference images. Early NR-IQA methods, such as NIQE(NIQE), primarily rely on hand-crafted Natural Scene Statistics (NSS) features. With the advancement of deep learning, many works have turned to neural networks for NR-IQA. For instance, MetaIQA(MetaIQA)leverages meta-learning to enhance model generalization. Additionally, methods such as CLIP-IQA(CLIP-IQA)leverage text prompts to enhance perceptual capabilities(sheng2023aesclip; yang2025language; li2025cgiaa). Given the powerful learning capabilities of large multi-modality models (LMMs)(sheng2025instructcrop; sheng2025agiqa), emerging work, such as Q-Align(Q-align), has begun to apply these models to quality assessment. Most recently, methods represented by Co-Instruct(Co-Instruct)have extended LMMs to compare multiple images. Furthermore, VisualQuality-R1(wu2025visualquality) introduces a novel methodology for IQA that leverages LMMs within a reinforcement learning-to-rank framework.
Although these datasets and methods have contributed significantly to the IQA community, most methods are score-based and cannot effectively handle fine-grained perceptual differences brought by image restoration algorithms. Recent works have explored fine-grained quality assessment in specific domains. Zhang et al. (zhang2019fine; zhang2022fine) introduced fine-grained assessment for compressed images, focusing on distinguishing compression artifacts within similar quality levels. However, these prior works primarily target image compression, whereas our work specifically addresses the unique challenges of fine-grained quality assessment in image restoration tasks, where the evaluation focus shifts from compression artifacts to perceptual restoration quality differences.
@@
In image restoration tasks, both algorithm comparison and optimization frequently involve evaluating images with subtle quality differences. Algorithm comparison requires distinguishing between restoration results with marginal quality differences, while parameter optimization involves incremental quality changes that demand sensitive assessment methods to identify optimal configurations. To investigate whether existing IQA methods can objectively capture fine-grained quality differences in IR task, we conduct a comprehensive computational analysis on established IQA datasets. Specifically, we evaluate state-of-the-art IQA methods to assess their fine-grained discrimination capabilities. Our evaluation encompasses both full-reference (FR) and no-reference (NR) methods. FR methods include PSNR, SSIM (SSIM), LPIPS (LPIPS), and DISTS (DISTS). NR methods include traditional approaches such as NIQE (NIQE), IL-NIQE (IL-NIQE), and BRISQUE (BRISQUE), as well as deep learning-based methods including DB-CNN (DBCNN), HyperIQA (HyperIQA), MetaIQA (MetaIQA), LIQE (LIQE), CLIP-IQA (CLIP-IQA), Q-Align (Q-align), and DeQA-Score (DeQA).
Specifically, we partition the quality score range into several intervals to examine how IQA methods perform when evaluating images with similar quality levels, which is particularly relevant for fine-grained assessment scenarios in image restoration. Table 2 presents the performance comparison across different MOS ranges on the PIPAL dataset (PIPAL). While most IQA methods achieve reasonable overall SRCC values on the whole dataset, their performance dramatically deteriorates when evaluated within narrow quality ranges. We also conducted similar analyses on other IQA datasets, which consistently demonstrate the same pattern. More results are provided in the supplementary material. These findings provide strong evidence for the inadequacy of current score-based IQA approaches in fine-grained scenarios, motivating the development of our fine-grained evaluation framework for image restoration.
@@
We collect images from multiple restoration-specific datasets to cover different degradation and visual appearances: 1) deraining images from IVIPC-DQA (IVIPC_DQA); 2) deblurring images from MDD13 (MDD13); 3) dehazing images from IVC-Dehazing (IVC-Dehazing) and exBeDDE (exBeDDE); 4) denoising images from PIPAL (PIPAL) and KADID-10k (KADID-10k); 5) super-resolution images from QADS (QADS), RealSRQ (RealSRQ), SISRSet (SISRSet), and PIPAL (PIPAL); 6) mixture restoration images from PIPAL (PIPAL). After image collection, we generate image pairs by randomly pairing all images with identical content and restoration tasks, resulting in a total of 1,275,297 image pairs.
@@
To remove unnoticeable pairs, we propose a JND-based filtering approach. The JND represents the minimum perceptual threshold for noticeable quality differences, making it ideal for identifying imperceptible quality pairs. Our JND-based filtering process operates as follows: (1) randomly select 20,000 images from the collected datasets, (2) compute the JND map for each image using established JND estimation model (wu2017enhanced), and (3) generate JND noise images by adding these JND maps as noise to the original images:
@@
where $I$ is the original image and $\text{JND}(I)$ is the corresponding JND noise map. We calculate SSIM (SSIM) between each $I$ and $I_{\text{JND}}$, and use the median SSIM value $\text{SSIM}_{\text{med}}$ as our filtering threshold. The unnoticeable pair filtering criterion is then defined as:
@@
After applying both filtering steps, we retain 30,886 fine-grained image pairs and their corresponding 18,408 images, forming the core of our FGRestore dataset. Due to space limitations in the main paper, we are unable to clearly demonstrate fine-grained restoration image pair examples at sufficient resolution. Additional examples of fine-grained restoration image pairs are provided in the supplementary material.
@@
To validate the necessity of fine-grained pairwise ranking annotations, we conduct a consistency analysis between MOS-based scoring and human preference ranking. We create a scatter plot comparing normalized MOS differences ($x$-axis) with human preference annotations ($y$-axis). As shown in Figure 4, the scatter plot reveals significant inconsistencies between MOS scores and human preferences. Points in the red regions represent inconsistent annotations where MOS scores and ranking preferences disagree. Critically, image pairs with smaller MOS differences exhibit larger scatter points, indicating that inconsistencies become more severe as quality differences become subtler. This strongly supports the necessity of fine-grained pairwise ranking, as MOS-based scoring becomes unreliable for discriminating subtle quality differences.
@@
We compare the proposed FGResQ model against state-of-the-art FR and NR IQA methods. For fair comparison, all learning-based methods with public source codes are retrained using the same training-testing protocol with our scene-aware loss replacing their original regression loss, except for Compare2Score (Compare2Score) which operates in a zero-shot manner without training. We evaluate performance using SRCC and PLCC for quality score prediction, and accuracy (ACC) for fine-grained pairwise ranking. We split the dataset at an 8:2 ratio by image pairs. Individual image division following the corresponding pair.
@@
We use CLIP ViT-B/16 as the backbone encoder for both degradation-aware feature learning and dual-branch quality prediction. Training images are resized to $256\times 256$ pixels and randomly cropped to $224\times 224$ patches. The model is trained on NVIDIA RTX4090 GPUs using the Adam optimizer with cosine annealing. The maximum learning rate is $5\times 10^{-6}$ and the batch size is 64. To ensure training stability, we employ scene-aware sampling strategy that groups samples from the same source dataset within each batch. $\lambda_{1}$ and $\lambda_{2}$ are set to 5 and 1, respectively.

### v4

@@ Abstract
Recent years have witnessed remarkable achievements in perceptual image restoration (IR), creating an urgent demand for accurate image quality assessment (IQA), which is essential for both performance comparison and algorithm optimization. Unfortunately, the existing IQA metrics exhibit inherent weakness for IR task, particularly when distinguishing fine-grained quality differences among restored images. To address this dilemma, we contribute the first-of-its-kind fine-grained image quality assessment dataset for image restoration, termed FGRestore, comprising 18,408 restored images across six common IR tasks. Beyond conventional scalar quality scores, FGRestore was also annotated with 30,886 fine-grained pairwise preferences. Based on FGRestore, a comprehensive benchmark was conducted on the existing IQA metrics, which reveal significant inconsistencies between score-based IQA evaluations and the fine-grained restoration quality. Motivated by these findings, we further propose FGResQ, a new IQA model specifically designed for image restoration, which features both coarse-grained score regression and fine-grained quality ranking. Extensive experiments and comparisons demonstrate that FGResQ significantly outperforms state-of-the-art IQA metrics.
Homepagehttps://sxfly99.github.io/FGResQ-Home/
@@ Introduction
Perceptual image restoration (IR) stands as a cornerstone in low-level computer vision, aiming to recover high-quality images from their degraded observations while maintaining perceptual fidelity (Potlapalli et al. 2023; Luo et al. 2023). Recent years have witnessed remarkable breakthroughs in this field, largely driven by the evolution of generative models. These advances have enabled IR algorithms to achieve unprecedented visual quality, creating an urgent demand for accurate image quality assessment (IQA) methods that can reliably evaluate and compare restored images. Such assessment capabilities are essential not only for performance benchmarking across different IR algorithms but also for guiding algorithmic optimization.
@@
Currently, IR algorithm evaluation still predominantly relies on traditional reference-based metrics such as PSNR and SSIM (Wang et al. 2004), which assess restoration quality by measuring the similarity between the restored images and ground-truth images, even though their inconsistency with human perceptual judgment has been increasingly observed by the research community (Jinjin et al. 2020). Recent efforts have introduced no-reference IQA methods (Xu et al. 2024; Zhou et al. 2025), including CLIP-IQA (Wang et al. 2023) and other learning-based approaches (Ke et al. 2021), which prove particularly valuable in real-world restoration scenarios where ground truth images are unavailable. However, a critical oversight persists: these methods have been primarily designed for image quality assessment, without specific consideration for the fine-grained evaluation requirements inherent in IR tasks.
@@
To investigate this question, we first conducted comprehensive computational analyses on established IQA datasets for image restoration (Jinjin et al. 2020; Shi et al. 2019). Our findings consistently reveal a significant performance gap: while existing IQA methods achieve reasonable results in coarse-grained quality assessment, they consistently underperform in fine-grained scenarios, with significant inconsistencies between predicted scores and human perceptual judgments. This observation highlights the inadequacy of current evaluation protocols for capturing the nuanced quality differences which are crucial in IR applications. Moreover, when quality differences between images become subtle, both human and IQA models encounter significant challenges in providing reliable absolute quality scores. In contrast, pairwise comparison(Yang et al. 2024b, a) presents a more reliable alternative for fine-grained assessment, as humans demonstrate superior consistency in relative quality judgments compared to absolute scoring.
@@
To evaluate IR algorithms, numerous datasets have been proposed. These include datasets dedicated to single restoration types, such as the MDD13 (Liu et al. 2013) dataset for deblurring, the IVC-Dehazing (Ma et al. 2015) dataset for dehazing, the IVIPC-DQA (Wu et al. 2019) dataset for deraining, and the QADS (Zhou et al. 2019) and SRIQA-Bench (Chen et al. 2025) datasets for super-resolution. Additionally, there are datasets that encompass distortions from various algorithms, such as PIPAL (Jinjin et al. 2020), which includes distortion types like denoising, super-resolution, and super-resolution with post-denoising, KADID-10k (Lin et al. 2019), which contains various image distortions including denoising degradations and DiffIQA(Chen et al. 2025), which accounts for degradation artifacts induced by diffusion-based image enhancement. A brief comparison of FGRestore and previous IR evaluation datasets is illustrated in Table 1.
@@
Existing image restoration (IR) algorithms typically employ full-reference image quality assessment (FR-IQA) metrics for performance evaluation, such as the conventional PSNR and SSIM (Wang et al. 2004), as well as deep learning-based metrics like LPIPS (Zhang et al. 2018a) and DISTS (Ding et al. 2020). To overcome the flawed ”perfect reference” assumption in traditional FR-IQA, which is challenged by imaging system limitations and superior generative methods, A-FINE (Chen et al. 2025) proposes a generalized model that adaptively assesses both the fidelity and naturalness of test images.
No-reference IQA (NR-IQA) methods have been proposed to assess image quality without reference images. Early NR-IQA methods, such as NIQE(Mittal et al. 2012b), primarily rely on hand-crafted Natural Scene Statistics (NSS) features. With the advancement of deep learning, many works have turned to neural networks for NR-IQA. For instance, MetaIQA(Zhu et al. 2020)leverages meta-learning to enhance model generalization. Additionally, methods such as CLIP-IQA(Wang et al. 2023)leverage text prompts to enhance perceptual capabilities(Sheng et al. 2023; Yang et al. 2025; Li et al. 2025). Given the powerful learning capabilities of large multi-modality models (LMMs)(Sheng et al. 2025a, b), emerging work, such as Q-Align(Wu et al. 2023), has begun to apply these models to quality assessment. Most recently, methods represented by Co-Instruct(Wu et al. 2024)have extended LMMs to compare multiple images. Furthermore, VisualQuality-R1(Wu et al. 2025) introduces a novel methodology for IQA that leverages LMMs within a reinforcement learning-to-rank framework.
Although these datasets and methods have contributed significantly to the IQA community, most methods are score-based and cannot effectively handle fine-grained perceptual differences brought by image restoration algorithms. Recent works have explored fine-grained quality assessment in specific domains. Zhang et al. (Zhang et al. 2018b, 2021) introduced fine-grained assessment for compressed images, focusing on distinguishing compression artifacts within similar quality levels. However, these prior works primarily target image compression, whereas our work specifically addresses the unique challenges of fine-grained quality assessment in image restoration tasks, where the evaluation focus shifts from compression artifacts to perceptual restoration quality differences.
@@
In image restoration tasks, both algorithm comparison and optimization frequently involve evaluating images with subtle quality differences. Algorithm comparison requires distinguishing between restoration results with marginal quality differences, while parameter optimization involves incremental quality changes that demand sensitive assessment methods to identify optimal configurations. To investigate whether existing IQA methods can objectively capture fine-grained quality differences in IR task, we conduct a comprehensive computational analysis on established IQA datasets. Specifically, we evaluate state-of-the-art IQA methods to assess their fine-grained discrimination capabilities. Our evaluation encompasses both full-reference (FR) and no-reference (NR) methods. FR methods include PSNR, SSIM (Wang et al. 2004), LPIPS (Zhang et al. 2018a), and DISTS (Ding et al. 2020). NR methods include traditional approaches such as NIQE (Mittal et al. 2012b), IL-NIQE (Zhang et al. 2015), and BRISQUE (Mittal et al. 2012a), as well as deep learning-based methods including DB-CNN (Zhang et al. 2020), HyperIQA (Su et al. 2020), MetaIQA (Zhu et al. 2020), LIQE (Zhang et al. 2023), CLIP-IQA (Wang et al. 2023), Q-Align (Wu et al. 2023), and DeQA-Score (You et al. 2025).
Specifically, we partition the quality score range into several intervals to examine how IQA methods perform when evaluating images with similar quality levels, which is particularly relevant for fine-grained assessment scenarios in image restoration. Table 2 presents the performance comparison across different MOS ranges on the PIPAL dataset (Jinjin et al. 2020). While most IQA methods achieve reasonable overall SRCC values on the whole dataset, their performance dramatically deteriorates when evaluated within narrow quality ranges. We also conducted similar analyses on other IQA datasets, which consistently demonstrate the same pattern. More results are provided in the supplementary material. These findings provide strong evidence for the inadequacy of current score-based IQA approaches in fine-grained scenarios, motivating the development of our fine-grained evaluation framework for image restoration.
@@
We collect images from multiple restoration-specific datasets to cover different degradation and visual appearances: 1) deraining images from IVIPC-DQA (Wu et al. 2019); 2) deblurring images from MDD13 (Liu et al. 2013); 3) dehazing images from IVC-Dehazing (Ma et al. 2015) and exBeDDE (Zhao et al. 2020); 4) denoising images from PIPAL (Jinjin et al. 2020) and KADID-10k (Lin et al. 2019); 5) super-resolution images from QADS (Zhou et al. 2019), RealSRQ (Jiang et al. 2022), SISRSet (Shi et al. 2019), and PIPAL (Jinjin et al. 2020); 6) mixture restoration images from PIPAL (Jinjin et al. 2020). After image collection, we generate image pairs by randomly pairing all images with identical content and restoration tasks, resulting in a total of 1,275,297 image pairs.
@@
To remove unnoticeable pairs, we propose a JND-based filtering approach. The JND represents the minimum perceptual threshold for noticeable quality differences, making it ideal for identifying imperceptible quality pairs. Our JND-based filtering process operates as follows: (1) randomly select 20,000 images from the collected datasets, (2) compute the JND map for each image using established JND estimation model (Wu et al. 2017), and (3) generate JND noise images by adding these JND maps as noise to the original images:
@@
where $I$ is the original image and $\text{JND}(I)$ is the corresponding JND noise map. We calculate SSIM (Wang et al. 2004) between each $I$ and $I_{\text{JND}}$, and use the median SSIM value $\text{SSIM}_{\text{med}}$ as our filtering threshold. The unnoticeable pair filtering criterion is then defined as:
@@
After applying both filtering steps, we retain 30,886 fine-grained image pairs and their corresponding 18,408 images, forming the core of our FGRestore dataset. Due to space limitations, high-resolution fine-grained restoration examples are provided in the supplementary material.
@@
To validate the necessity of fine-grained pairwise ranking, we analyze the consistency between MOS-based scoring and human preferences. Figure 4 compares normalized MOS differences ($x$-axis) with human preferences ($y$-axis), revealing significant inconsistencies. Points in the red regions represent cases where MOS scores and preferences disagree. Critically, image pairs with smaller MOS differences exhibit larger scatter, indicating that inconsistencies worsen as quality differences become subtler. This supports the necessity of fine-grained pairwise ranking, as MOS scoring becomes unreliable for discriminating subtle quality differences.
@@
We compare the proposed FGResQ model against state-of-the-art FR and NR IQA methods. For fair comparison, all learning-based methods with public source codes are retrained using the same training-testing protocol with our scene-aware loss replacing their original regression loss, except for Compare2Score (Zhu et al. 2024) which operates in a zero-shot manner without training. We evaluate performance using SRCC and PLCC for quality score prediction, and accuracy (ACC) for fine-grained pairwise ranking. We split the dataset at an 8:2 ratio by image pairs. Individual image division following the corresponding pair.
@@
We use CLIP ViT-B/16 as the backbone encoder for degradation-aware feature learning and dual-branch quality prediction. The model is trained on NVIDIA RTX4090 GPUs using the Adam optimizer with cosine annealing. The maximum learning rate is $5\times 10^{-6}$ and the batch size is 64. To ensure training stability, we employ scene-aware sampling that groups samples from the same source dataset within each batch. $\lambda_{1}$ and $\lambda_{2}$ are set to 5 and 1, respectively.